# Dictionary-Learning-Based Data Pruning for System Identification

### v1

@@ 3.2 Data with Dual Stable Equilibria
Figure 5: Phase portrait of the nonlinear system with dual stable equilibria
It can be seen that this system exhibits two stable equilibria. To comprehensively capture its dynamics, the data pruning algorithm should identify two distinct types of samples attracted to the different stable spirals and select samples from both categories.
Unlike feature selection, random selection serves as a robust benchmark for sample selection and often outperforms more complex algorithms [26]. Therefore, the results of the proposed data-pruning method will be compared with those of the random selection method to demonstrate its advantages.
The structure of the dual-stable-equilibria (DSE) data is visualized using its first and second principal components. The selected samples and atoms are projected onto these principal directions, with their distribution shown in Fig. 6. It can be found that the samples selected via mini-batch sample selection are distributed across the entire dataset, whereas the randomly selected samples are concentrated primarily on the left and right regions, with less coverage in the middle.
Figure 6: Visualization of the DSE data structure
@@
To quantitatively evaluate the selection performance of the mini-batch and random methods, a reduced polynomial NARX model with 10 terms is derived using the full dataset as the baseline. Note that in all case studies, the number of terms for system output and input is ten respectively, so the total number of terms is $20$, and the degree of polynomial nonlinearity is three. The model coefficients trained on the full dataset are used as the baseline results. The performance of the baseline NARX on the test dataset is provided in Appendix 1. The NARX terms are then fixed, and the model coefficients are trained using the pruned datasets obtained via the mini-batch and random selection methods. The number of atoms for dictionary learning is set to 100, and 600 samples are then selected. The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared. The sample selection and training processes are repeated ten times, with the results presented in Fig. 7.
The mini-batch method gives results with a higher median and lower variance, whereas the random selection method performs significantly worse. The same phenomenon is observed when the number of selected samples varies between 300 and 600, as illustrated in Fig. 8. Furthermore, for the mini-batch method, the variance of the model coefficients decreases as the number of selected samples increases. However, this relationship does not exist when the samples are randomly selected.
@@
Figure 7: Comparison of sample selection performance between mini-batch and random selection methods on DSE data using R-squared
@@
Figure 8: The effect of sample size on the performance of mini-batch and random selection methods for DSE data
@@ 4 Case Studies on the Benchmark Datasets
Two benchmark datasets, collected from the Electro-Mechanical Positioning System (EMPS) [27] and the Wiener-Hammerstein System (WHS) [28] are adopted to evaluate the performance of the mini-batch method for data pruning in real system identification.
@@ 4.1 Data from the Electro-Mechanical Positioning System
The structure of the data collected from the EMPS, a standard drive configuration for prismatic joints of robots or machine tools, is visualised by its first two principal components, as shown in Fig. 9. The primary nonlinearity of the corresponding data is introduced by friction effects.
The atoms determined for the dictionary learning and the selected samples are also shown in Fig 9. The samples selected by the mini-batch method are spread over all parts of the dataset, while the randomly selected samples are concentrated in the middle two parts but pay less attention to the left and the right parts.
Figure 9: Visualization of the EMPS data structure
Figure 10: Comparison of sample selection performance between mini-batch and random selection methods on EMPS data using R-squared
The baseline NARX model with 10 terms is derived again, and the model coefficients trained on the full dataset are used as the baseline to quantitatively evaluate the selection performance of the mini-batch and random methods. See Appendix 1 for the performance of the baseline NARX on the test dataset. The number of atoms for dictionary learning is set to 100, and 10000 samples are then selected. The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared, with the NARX terms fixed. The sample selection and training processes are repeated ten times, with the results illustrated in Fig. 10. In addition, Fig. 11 shows the results corresponding to two methods when the number of selected samples varies between 2000 and 20000.
Figure 11: The effect of sample size on the performance of mini-batch and random selection methods for EMPS data
Similar to the observations from the DSE data, the mini-batch method generally produces results with a higher median and lower variance. However, there are exceptions when the number of selected samples is less than 6000, where the randomly selected samples yield higher mean values. This scenario will be analysed and further discussed in Section 4.3. Additionally, as expected, the variance of the model coefficients decreases as the number of selected samples increases.
@@ 4.2 Data from the Wiener-Hammerstein System
The structure of the data collected from the Wiener-Hammerstein System (WHS), a well-known block-oriented structure consisting of a static nonlinearity between two linear time-invariant blocks, is visualised in Fig. 12 using its first two principal components. The atoms determined for the dictionary learning and the selected samples are also shown in Fig 14. Different from the observations from the DSE data and the EMPS data, the samples obtained by the mini-batch method and random selection method have similar distributions, with the selected samples distributed over the entire dataset.
@@
The same strategy is applied to quantitatively evaluate the sample selection performance of the mini-batch and random methods for the identification task of WHS. The baseline NARX model with 10 terms is derived, and its performance on the test dataset is provided in Appendix 1. The number of atoms for dictionary learning is set to 100, and 10000 samples are selected, with the R-squared results shown in Fig. 13. Then, the number of selected samples varies between 2000 and 20000 to demonstrate the variation of R-squared with the selected sample size, as presented in Fig. 14.
Figure 12: Visualization of the WHS data structure
Figure 13: Comparison of sample selection performance between mini-batch and random selection methods on WHS data using R-squared
Figure 14: The effect of sample size on the performance of mini-batch and random selection methods for WHS data
Fig. 13 demonstrates that the results obtained by the mini-batch and random methods are similar. This phenomenon is also observed in Fig. 14 when the number of the selected samples exceeds 8000. This result is reasonable and expected, as similar sample distributions provide similar information for system identification. In addition, similar to the EMPS case, when the number of selected samples is small (less than 8000), the randomly selected samples yield higher mean values. This phenomenon will be analysed and further discussed in the following section.
4.3 The Effect of Hyperparameters in the Mini-batch Method
In the mini-batch method, sample selection results are influenced by two hyperparameters: the number of atoms in the dictionary and the batch size. To explore how these two hyperparameters affect the performance of the selected samples, system identification results against the varying atom size and batch size for the three previously used datasets are given in Fig. 15 to Fig. 16. For the DSE dataset, $600$ samples are selected, while for EMPS and WHS datasets, $10000$ samples are selected.
It can be noticed that the batch size in Algorithm 1 is determined by both the number of samples and the number of atoms, with the aim of minimising redundancy among the selected samples when all atoms are used in the sample selection process. Therefore, for the results in Fig. 15, the batch size varies for each atom size. Notably, a larger atom size does not leads to better performance, because some atoms may be learned from noise rather than meaningful data. In Fig. 16, when changing the batch size for each dataset, the number of atoms is fixed at the best atom size obtained from Fig. 15. In this scenario, using a larger batch size to reduce redundancy can improve the performance of the selected samples.
(a) DSE dataset
(b) EMPS dataset
(c) WHS dataset
@@ 4.3 The Effect of Hyperparameters in the Mini-batch Method
Figure 15: The effect of atom size on the performance of mini-batch method
(a) DSE dataset
(b) EMPS dataset
(c) WHS dataset
@@
Figure 16: The effect of batch size on the performance of mini-batch method with the best atom size
To investigate whether a larger batch size always leads to better selection when the number of atoms is fixed, the relationship between R-squared and batch size for EMPS data with different atom sizes is presented in Fig. 17. The results indicate that the relationship between R-squared and batch size is not a straightforward linear relationship.
In summary, the optimal atom size is different for each dataset, and the optimal batch size may not be the ratio of the sample number to the atom size. A more sophisticated method is required to determine the appropriate hyperparameters for the proposed method. Furthermore, the possible reason for the previous-mentioned phenomenon shown in Fig. 11 and Fig. 14 is that when the number of selected samples is small, the performance of the selected samples may be more sensitive to the effectiveness of the hyperparameters.
(a) atom size 10
(b) atom size 100
(c) atom size 1000
@@
Figure 17: The effect of batch size on the performance of mini-batch method with difference atom sizes for EMPS data
5 Conclusion
This paper introduces a data-pruning method based on dictionary learning to choose appropriate time-series samples for identification tasks of a discrete system with finite orders. The selection process is independent of the specific system identification technique used.
@@ 5 Conclusion
The case studies on the synthetic and benchmark datasets show that the proposed method effectively reduces the sample size without significantly degrading identification performance. Compared with random selection, the data-pruning method gives better results with a higher mean and lower variance of R-squared values. Additionally, it is found that the careful selection of hyperparamters is crucial, particularly when the number of selected samples is small.
@@
Figure 18: The performance of the baseline NARX model for DSE test data with different initial conditions
Figure 19: The performance of the baseline NARX model for EMPS test data
Figure 20: The performance of the baseline NARX model for WHS test data

### v2

@@ 3.2 Data with Dual Stable Equilibria
Figure 4: A phase portrait of the nonlinear system with dual stable equilibria
It can be seen that this system exhibits two stable equilibria, with five measurements for each of the left and right equilibria. Therefore, this dataset is referred to as symmetrical dual-stable-equilibria (SDSE) dataset. To comprehensively capture its dynamics, the data pruning algorithm should identify two distinct kinds of samples attracted to the different stable spirals and select samples from both kinds. Unlike feature selection, random selection serves as a strong baseline for sample selection and often outperforms more sophisticated algorithms when retaining 30% or less of the data [25]. Therefore, the results of the proposed data-pruning method will be compared with those of the random selection method to demonstrate its advantages.
@@
To quantitatively evaluate the selection performance of the mini-batch FastCan and random methods, a reduced polynomial NARX model with ten terms and an intercept is derived using the full dataset as the baseline. The model coefficients trained on the full training dataset are used as the baseline results. A more detailed description of the baseline NARX model and its prediction performance on the test dataset are provided in Appendix 1. The NARX terms are then fixed, and the model coefficients are trained using the pruned training datasets obtained via the mini-batch FastCan and random selection methods. 100 samples are then selected as an example, and the corresponding number of atoms for dictionary learning is set to 15. The reasons for this number of atoms will be discussed in Section 5. The coefficients trained on the two pruned datasets are compared with those trained on the full dataset using the R-squared. The sample selection and training processes are repeated ten times, with the results presented in Fig. 5a.
(a) SDSE dataset
(b) ADSE dataset
@@
Figure 5: Comparison of sample selection performance between mini-batch FastCan and random selection methods using R-squared on two datasets
As shown in Fig. 5a, the mini-batch FastCan method gives results with a higher median and lower variance, indicating superior and more consistent performance compared to random selection. As illustrated in Fig. 6a, the same pattern is observed when the number of selected samples varies between 50 and 150 and the number of atoms is fixed to 15, except for 60 samples—this case will be discussed in Section 5. Furthermore, for the mini-batch FastCan method, the variance of the model coefficients decreases as the number of selected samples increases. However, this relationship does not exist when the samples are randomly selected in Fig. 6a.
(a) SDSE dataset
(b) ADSE dataset
@@
Figure 6: The effect of sample size on the performance of mini-batch FastCan and random selection methods for two datasets. SD stands for standard deviation.
To examine whether the sample distributions corresponding to the results shown in Fig. 5a are different, the structure of the SDSE data is visualised using its first and second principal components, as shown in Fig. 7a. The selected samples corresponding to the results shown in Fig. 5a are projected onto these principal directions, with their distribution shown in Fig. 7a. It can be observed that the samples selected via mini-batch FastCan tend to cluster around the learned atoms, and the randomly selected samples are evenly distributed across the entire dataset according to the spatial distribution of candidate samples. Compared to random selection, the proposed method makes more samples distributed at the left and right ends. Nevertheless, the overall difference between the two sample distributions is relatively small.
(a) SDSE dataset
(b) ADSE dataset
Figure 7: Visualisation of the data structure of two datasets
To further determine whether the sample distribution selected by the proposed method is different from that of the random method, 2 measurements in left equilibrium and 98 measurements in right equilibrium are simulated to repeat the previous analysis. This dataset is referred to as asymmetrical dual-stable-equilibria (ADSE) dataset. 100 samples are selected, and the number of atoms is set to 20 (see Section 5 for details on this setting). The results are presented in Figs. 5b-7b. As shown in Figs. 5b and 6b, when the number of samples across different conditions is imbalanced, the mini-batch FastCan method has obvious advantages over the random sample selection method. Furthermore, as shown in Fig. 7b, almost all samples selected by the random method are concentrated in the right area, that is, these samples belong to the right equilibrium. In contrast, the proposed method selects samples from both equilibria, with those from the left equilibrium evenly distributed in the left area. These phenomena indicate that, unlike random selection, the mini-batch FastCan method is less influenced by the imbalance in the candidate sample distribution and is more effective at capturing diverse system behaviours. This phenomenon, in turn, helps explain why the identification performance corresponding to the proposed method is superior to that corresponding to the random selection method.
@@ 4 Case Studies on the Benchmark Datasets
Two benchmark datasets, collected from the Electro-Mechanical Positioning System (EMPS) [26] and the Wiener-Hammerstein System (WHS) [27] are adopted to evaluate the performance of the mini-batch FastCan method for data pruning in real-world system identification tasks.
@@ 4.1 Data from the Electro-Mechanical Positioning System
The EMPS is a standard drive configuration for prismatic joints of robots or machine tools. The primary nonlinearity of the corresponding data is introduced by friction effects. The baseline NARX model with ten terms and an intercept is derived, and the model coefficients trained on the full training dataset are used as the baseline to quantitatively evaluate the selection performance of the mini-batch FastCan and random methods. See Appendix 1 for the performance of the baseline NARX on the test dataset.
The number of selected samples is 100, and the number of atoms for dictionary learning is set to 25 (see Section 5 for details on this setting). The coefficients trained on the two pruned datasets are compared with those from the full dataset using the R-squared, with the NARX terms fixed. The sample selection and training processes are repeated ten times, with the results illustrated in Fig. 8a. In addition, Fig. 9a presents the results corresponding to two methods for different numbers of selected samples, ranging from 20 to 120.
(a) EMPS dataset
(b) WHS dataset
Figure 8: Comparison of sample selection performance between mini-batch FastCan and random selection methods using R-squared on two benchmark datasets
(a) EMPS dataset
(b) WHS dataset
Figure 9: The effect of sample size on the performance of mini-batch FastCan and random selection methods for two benchmark datasets. SD stands for standard deviation.
Consistent with the observations from the SDSE and ADSE datasets, the mini-batch FastCan method produces results with a higher median and lower variance, indicating more stable and reliable performance. Additionally, as expected, the variance of the model coefficients also decreases as the number of selected samples increases for the mini-batch FastCan method.
The structure of the data collected from the EMPS is visualised by its first two principal components, as shown in Fig. 10a. The atoms determined for the dictionary learning and the selected samples corresponding to the results given in Fig. 8a are projected onto this reduced space. The atoms and selected samples corresponding to the mini-batch FastCan method are distributed in the central, left and right regions, while the randomly selected samples are almost only spread over the central part.
(a) EMPS dataset
(b) WHS dataset
Figure 10: Visualisation of the data structure of two benchmark datasets
@@ 4.2 Data from the Wiener-Hammerstein System
The WHS is a well-known block-oriented structure consisting of a static nonlinearity between two linear time-invariant blocks. The same strategy is applied to quantitatively evaluate the sample selection performance of the mini-batch FastCan and random methods for the identification task of WHS. The baseline NARX model with ten terms and an intercept is derived, and its performance on the test dataset is provided in Appendix 1.
The number of atoms for dictionary learning is set to 5 (see Section 5 for details on this setting, and 100 samples are selected, with the R-squared results shown in Fig. 8b. Then, the number of selected samples varies between 20 and 120 to demonstrate the variation of R-squared with the selected sample size, as presented in Fig. 9b. Fig. 8b demonstrates that the results obtained by the mini-batch FastCan are better than those obtained by the random method. This phenomenon is also observed in Fig. 9b except at 40 samples. This exception will be analysed and further discussed in the following section.
The structure of the data collected from the WHS is visualised in Fig. 10b using its first two principal components. The atoms determined for the dictionary learning and the selected samples are shown in Fig. 10b, with the number of atoms set to 5. Different from the observations from the ADSE data and the EMPS data, the samples obtained by the mini-batch FastCan method and the random selection method have similar distributions, with the selected samples mainly distributed over the central part of the dataset. This phenomenon is unexpected because the samples associated with the results in Fig. 8b exhibit similar distributions but provide varying levels of information for system identification. This observation suggests that sample selection strategies based solely on distributional characteristics of the data space may be insufficient for identifying informative samples.
(a) SDSE dataset
(b) ADSE dataset
(c) EMPS dataset
(d) WHS dataset
@@
Figure 11: The effect of atom size on the performance of mini-batch FastCan method
5 The Effect of Hyperparameters in the Mini-batch FastCan Method
In the mini-batch FastCan method, sample selection results are influenced by two hyperparameters: the number of atoms in the dictionary and the batch size. To examine how these two hyperparameters affect the performance of the selected samples, system identification results against the varying atom size and batch size for the four previously used datasets are given in Fig. 11 to Fig. 12. For all datasets, $100$ samples are selected. For the results in Fig. 11, the batch size is not given as an input parameter but is instead set to the ceiling of the ratio between the number of selected samples and the number of atoms ($\lceil n/q\rceil$), with the aim of minimising redundancy among the selected samples when all atoms are used in the sample selection process. Therefore, the batch size varies for each atom size in Fig. 11.
As shown in Fig. 11, the performance deteriorates when the atom size exceeds the size associated with the highest R-squared value. This may be attributed to the fact that, for a fixed number of selected samples, larger atom sizes lead to smaller batch sizes, thus increasing sample-wise redundancy. In practical engineering applications, the optimal atom size can be determined by defining a candidate range and selecting the value that yields the best performance via an exhaustive search.
(a) SDSE dataset
(b) ADSE dataset
(c) EMPS dataset
(d) WHS dataset
@@ 5 The Effect of Hyperparameters in the Mini-batch FastCan Method
Figure 12: The effect of batch size on the performance of mini-batch FastCan method under the optimal atom size
(a) SDSE dataset
(b) ADSE dataset
(c) EMPS dataset
(d) WHS dataset
@@
Figure 13: The effect of atom size on the performance of mini-batch FastCan method under different sample numbers
(a) SDSE dataset
(b) WHS dataset
Figure 14: Comparison of sample selection performance between the mini-batch FastCan method with tuned atom size and random selection using R-squared on two datasets
In Fig. 12, when changing the batch size for each dataset, the number of atoms is fixed at the optimal atom size obtained from Fig. 11, where the optimal atom sizes $q$ for the SDSE, ADSE, EMPS, and WHS datasets are 15, 20, 25, and 5, respectively. According to Step 2 in Algorithm 1, the maximum batch sizes for the DSE, EMPS, and WHS datasets are 7, 5, 4 and 10, respectively. It can be seen that when the atom size is fixed at the optimal value, using a larger batch size to reduce redundancy can generally improve the performance of the selected samples. Therefore, the atom sizes and the batch sizes for the previous four case studies are the optimal atom sizes obtained from Fig. 11 and the corresponding maximum batch sizes. However, it is worth noting that in Fig. 12, the relationship between R-squared and batch size is not a straightforward linear relationship, especially for the EMPS and WHS datasets. A more sophisticated method can be developed in the future to determine the appropriate hyperparameters for the proposed method.
To examine the underlying causes of the previously observed exceptions—where random selection outperforms or matches the performance of the mini-batch FastCan method, as shown in Figs. 6a and 9b—Fig. 13 presents the impact of atom size on model performance across varying selected-sample sizes for the mini-batch FastCan approach. This analysis aims to assess whether suboptimal hyperparameter settings contribute to these discrepancies.
To investigate the potential causes for the previously noted exceptions where random selection outperforms or matches mini-batch FastCan selection in Figs. 6a and 9b, the effect of atom size on model performance corresponding to the mini-batch FastCan method across different selected-sample sizes is demonstrated in Fig. 13. For each sample size, the maximum R-squared value is printed out. It can be observed that when the mini-batch FastCan method performs close to or worse than the random method, the adopted atom size is not the optimal size for the given number of selected samples. For the ADSE dataset, when 60 samples are selected, the optimal atom size is 5 rather than 15. For the WHS dataset, when 40 samples are selected, the optimal atom size is 10 rather than 5. After tuning the atom size for the mini-batch FastCan method, the comparison of sample selection performance of the two methods is then illustrated in Fig. 14. The results indicate that the mini-batch FastCan method can provide better sample selection results when the appropriate atom size is applied.
In addition, as shown in Figs. 13a, 13c and 13d, when the number of selected samples exceeds a certain threshold, the optimal atom size no longer changes with the number of selected samples. At the same time, Figs. 6a, 9a and 9b indicate that the performance of the selected samples also stabilises when the number of selected samples exceeds this threshold for each dataset. This consistent phenomenon suggests that when more samples cannot contribute additional information, the number of optimal atoms learned from the selected samples tends to stabilise. This finding supports, to a certain extent, the rationality and robustness of using dictionary learning to generate pseudo labels. Moreover, careful tuning of hyperparameters is particularly important when the number of selected samples is small.
6 Conclusions
This paper introduces a dictionary learning-based data pruning method, called mini-batch FastCan, for selecting informative time-series samples in the identification of discrete systems with finite orders. Specifically, k-means-based dictionary learning is well-suited for generating pseudo-labels in the data pruning process, mitigating the impact of imbalanced candidate sample distributions on sample selection. Additionally, selecting samples in separate batches allows for the selection of more samples than the number of necessary model terms, while effectively accounting for redundancy within each batch. This method helps ensure that the selected samples are more diverse and informative.
The selection process is independent of the specific system identification technique used, which enhances its versatility. Although hyperparameters, including atom size and batch size, can be empirically determined by engineering practices, the development of theoretically grounded optimisation strategies is recommended to improve generalizability and methodological rigour in future work.
@@ 6 Conclusions
The case studies on the synthetic and benchmark datasets show that the proposed method effectively reduces the sample size without significantly degrading the identification performance of the benchmark system. Compared to random selection, the data-pruning method generally yields better results, with a higher mean and lower variance of R-squared values. Additionally, it is found that the careful selection of hyperparameters is crucial, particularly when the proportion of selected samples is small.
Acknowledgements
The authors would like to acknowledge the support for this work from Shanghai Qi Zhi Institute Innovation Program (SQZ202310) and National Natural Science Foundation of China (52378187).
For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any arising Author Accepted Manuscript version.
@@ Acknowledgements
For each of the four datasets, a reduced polynomial NARX model with ten terms and an intercept is derived using the full dataset as a baseline. To select relevant nonlinear terms, the maximum input and output lags are set to 4 for the SDSE system, ADSE system, and EMPS, and to 7 for the WHS. For all systems, the polynomial degree is fixed at three, and the number of selected model terms is limited to ten. These parameter settings are designed to provide reasonable predictive performance, but can be further optimised based on specific application requirements.
@@
Figure 15: The performance of the baseline NARX model for SDSE test data with different initial conditions
(a) Test 1
(b) Test 2
Figure 16: The performance of the baseline NARX model for ADSE test data with different initial conditions
(a) EMPS dataset
(b) WHS dataset
Figure 17: The performance of the baseline NARX model for two benchmark datasets